OpenStudy (anonymous):

Solve showing all work. 3n - 5 = -8(6 + 5n)

OpenStudy (luigi0210):

Distribute the right side

OpenStudy (anonymous):

i forgot how to solve these type of problems :/

OpenStudy (luigi0210):

Well you know how to distribute right?

OpenStudy (anonymous):

i know that i have to add or subtract to both sides until its equal but i don't know what to add/subtract @Luigi0210

OpenStudy (luigi0210):

Well you need to DISTRIBUTE first..

OpenStudy (anonymous):

HOW?

OpenStudy (anonymous):

3n - 5 = -8(6 + 5n) 3n-5= -48 - 40n

OpenStudy (luigi0210):

Like that ^ Thanks you @hello1213

OpenStudy (anonymous):

but i don't understand how

OpenStudy (anonymous):

3n-5= -8(6) + -8(5n)

OpenStudy (anonymous):

okay.. what do i do after that?

OpenStudy (luigi0210):

Now you start moving things

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

what does that mean again? ugh sorry i totally forgot everything during summer break

OpenStudy (anonymous):

get the n numbers together on one side... the other numbers on the other side.

OpenStudy (anonymous):

the mission is to isolate n on one side... lets do the right side as the n numbers... what do you get?

OpenStudy (anonymous):

actually get the n numbers on the left side it is easier.

OpenStudy (anonymous):

i'm confused! can you show me how?

OpenStudy (anonymous):

i understand better by seeing how it's done rather than reading

OpenStudy (anonymous):

3n-5= -8(6) + -8(5n) 3n-5= -8(6) + -8(5n) <<< multiply 3n-5= -48 + (-40n) now isolate n numbers on left side 3n-5 + 40n= -48 - 40 n + 40n 3n-5+5+40n= -48 + 5 43n = -43

OpenStudy (anonymous):

Now what is the answer?

OpenStudy (anonymous):

-43?

OpenStudy (anonymous):

ohhh i think i get it now

OpenStudy (anonymous):

that is not the answer

OpenStudy (luigi0210):

@hello1213 try to avoid doing all of the asker's work for them

OpenStudy (anonymous):

nvm i don't get it -_-

OpenStudy (anonymous):

-43/43 is what

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

43n = -43 to isolate n divide BOTH sides by 43 get it?

OpenStudy (anonymous):

ohh yeahh! thanks :)

OpenStudy (anonymous):

No problem.